OpenStudy (anonymous):

Solve 2x – 5 = 3(2x + 5). You must show all work.

OpenStudy (anonymous):

First, work the parenthesis. 3(2x+5)= (3 x 2x) + (3 x 5). so, that makes it 6x + 15. Now the problem looks like this -2x - 5= 6x + 15. Next, move the variables to one side. Subtract both sides by 2x. (2x - 5) -2x = -5 (6x + 15) -2x = 4x + 15. The problem is now -5 = 4x + 15 Now, subtract 15 from both sides. -5 - 15 = -20. (4x + 15) - 15= 4x The problem is now -20 = 4x. All that's left is to simplify. Divide 4 to both sides. -20/4 = -5. 4x/4 = 1x. Your answer is -5 = x (x = -5).